OpenStudy (anonymous):

Finding the Limit Lim (x^2)-2x+3/(3x^2)+2x-5 x->2

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2} x ^{2}-2x+3/3x^{2}+2x-5\]

OpenStudy (anonymous):

do you get \(\frac{0}{0}\) if you substitute?

OpenStudy (anonymous):

if not, then whatever you get is the answer

OpenStudy (tkhunny):

Since what you have written is continuous at x = 2, it is simple: 4 - 4 + 1/4 + 4 - 5 = -3/4. Of course, if you meant \(\dfrac{x^{2} - 2x + 3}{3x^{2} + 2x - 5}\), that would give a different result.

OpenStudy (anonymous):

unless you get \(\frac{0}{a}\) and \(a\neq 0\) in which case there is no limit

OpenStudy (anonymous):

tkhunny, I meant the second problem you put

OpenStudy (anonymous):

a quick calculation gives \(\frac{3}{11}\) but you should check my arithmetic

OpenStudy (tkhunny):

@maink2 Well, then you should WRITE that. Please look at what you have written and take some good, careful time to reflect on the Order of Operations.

OpenStudy (anonymous):

That's what I got as well satellite73. Didn't know if there was more to it though